OpenStudy (anonymous):

how do i find the missing coordinate......y=-x-6 (3,y)

OpenStudy (anonymous):

plug in 3 for x. See what y equals.

OpenStudy (anonymous):

9

OpenStudy (anonymous):

\(-3-6 \ne 9\)

OpenStudy (anonymous):

y = -x - 6 = -(3) - 6 = ?

OpenStudy (saifoo.khan):

\[\Huge y = -9*\]